Solve the following linear equations. 
x/2 – 1/5 = x/3 + ¼ 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

x/2–1/5=x/3+1⁄4
⇒x/2–x/3=1⁄4+1/5
⇒(3x–2x)/6=(5+4)/20
⇒3x–2x=9/20×6
⇒x=54/20
⇒x=27/10

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
X/2 - 1/5 = X/3 + 1/4
X/2 - X/3 = 1/4 + 1/5
(3X - 2X)/6 = (5+4)/20
X/6 = 9/20
20 X = 54
X = 54/20 = 27/10
 
Hope it helps!
Thank you ✌️